One chocolate chip used in making chocolate chip cookies has a mass of 0.354 g. How many chocolate chips are there in one mole of chocolate chips?
1
Expert's answer
2016-09-19T07:47:18-0400
Elemental composition of chocolate is C4H8O4 with molar mass 120.1 g/mol. n(one chocolate chip) = 0.354 g/120.1 g/mol = 2.95*10-3 mol. Thus, number of chips is 1 mol/2.95*10-3 mol per chip = 339